OpenStudy (anonymous):

If r = -6, s = -2, t = -3 v = 10, and w = - 1 2 then wv|-t - s| = A) -25 B) -5 Eliminate C) 5 D) 25

OpenStudy (anonymous):

um its hard to read what you put but ill try LOL

OpenStudy (anonymous):

If r = -6, s = -2, t = -3 v = 10, and w = -1/2 then wv|-t - s| =

OpenStudy (anonymous):

this is quite easy just put the numbers with the letters sum them up to get your answer

OpenStudy (anonymous):

10x-0.5|3-2|= -17

OpenStudy (anonymous):

yeah but that is not a choice

OpenStudy (anonymous):

A) -25 B) -5 C) 5 Eliminate D) 25

OpenStudy (anonymous):

oh my bad i did it wrong hold on

OpenStudy (alexwee123):

A

OpenStudy (anonymous):

i got -25

OpenStudy (anonymous):

okay thanks you guys!